Massachusetts Institute of Technology Handout 0 6.042J/18.062J: Mathematics for Computer Science September 10, 1998 Professors Nancy Lynch and Srinivas Devadas

Information Sheet

Please return this sheet to us at the end of lecture today.

In addition to using this information to determine tutorial assignments, we will distribute a class list with phone numbers and email addresses to facilitate collaboration. If you do not want your phone number or email address distributed by us, please say so below. Name

Department and year

Phone number

Email address

Please list all Course VI and Course XVIII subjects you have taken or are currently taking. Circle those you are currently taking.

Each student must attend a two-hour tutorial on Friday. A list of sections is given below. Note that there are sixteen sections, however, some have been cancelled as marked below.

R01 F10-12 8-119            R02 F10-12 2-142           R03 F11-1 26-210
R04 F1-3  26-168            R05 F1-3  36-839           R06 F2-4  26-310
R07 F2-4  24-402            R08 F10-12 36-372          R09 F2-4  24-407
R10 F2-4  36-144            R11 9-11 26-168            R12 F3-5  26-168
R13 F11-1  2-139            R16 F11-1  2-147

You have already been assigned a tutorial by the Registrar. Please attend that tutorial, UNLESS it is cancelled or you have a difficult-to-change conflict. In that case, you may attend the recitation of your choice tomorrow. We will assign you to a permanent one by next week. Write the tutorial number you have been assigned by the Registrar:

If you have a conflict, write down the number of a tutorial that that you will attend tomorrow:

Regardless of whether you have a conflict or not, write down numbers of additional tutorials (if any) that you can attend starting next week: